### v2

# Structural Theory of Information Backflow in Non-Markovian Relaxation: TC/TCL Formalism and Minimal Phase Diagrams

Abstract
We develop a structural theory of information backflow in minimal non-Markovian relaxation processes within the framework of nonequilibrium statistical mechanics. The approach is based on the time-convolution (TC) and time-convolutionless (TCL) projection-operator formalisms for reduced dynamics and on the doubling construction of non-equilibrium thermo field dynamics, which provides an embedding representation of dissipative evolution.
We introduce a general backflow functional associated with a time-dependent information measure and derive generator-based sufficient conditions for the absence of backflow in terms of divisibility properties and effective relaxation rates. This allows a direct connection between memory kernels in generalized master equations and observable transient phenomena such as entropy overshoot and revival.
Furthermore, we propose a decomposition of backflow into classical mixing and intrinsic contributions in the doubled representation, leading to a unified classification of transient regimes. Minimal classical and quantum two-state models are analyzed as analytically tractable examples, yielding explicit phase diagrams and recovering Mittag-Leffler-type fractional relaxation as a universal envelope of non-Markovian damping.
The framework provides a constructive TC-to-TCL procedure for extracting effective rates and organizing memory-induced phenomena in a model-independent manner.
1 Introduction
The dynamics of nonequilibrium relaxation is often characterized by monotonic decay of information-theoretic quantities, such as entropies, relative entropies, and distinguishability measures. However, in the presence of memory effects, these quantities can exhibit non-monotonic transients: overshoots, revivals, and oscillatory relaxations. Such behaviors are commonly interpreted as information backflow from hidden degrees of freedom to the observable sector [10, 9, 8].
The goal of the present work is to elevate this interpretation from a qualitative narrative to a structural principle. We seek conditions under which a given information measure must relax monotonically, and conversely, structural mechanisms that necessarily permit backflow. A central technical tool is the projection-operator formalism, which yields both the time-convolution (TC) master equation and the time-convolutionless (TCL) master equation [1, 2, 4, 3, 6]. The TC form makes memory kernels explicit, whereas the TCL form yields a time-local generator that encodes the same memory effect. A second key ingredient is non-equilibrium thermo field dynamics (NETFD), in which density-operator evolution is represented as a Schrödinger-like evolution in a doubled Hilbert space. This provides a natural embedding interpretation and connects nonequilibrium relaxation to thermo-field entanglement.
The novelty of our approach is to combine TC/TCL structure with NETFD correspondence principles to derive general monotonicity theorems for information measures and to define a decomposition of backflow into classical and intrinsic entanglement sectors. This leads to a unified classification of transient phenomena, including entropy overshoots in classical stochastic models and entanglement revivals in minimal quantum systems.
Contributions
The main contributions of the present paper are summarized as follows:
  • We formulate information backflow as a universal functional $N_{I}$ in the TC/TCL reduced-dynamics framework, and connect it to divisibility properties of the TCL generator.
  • We provide generator-based sufficient conditions for the absence of backflow in both quantum (CP-divisible) and classical (stochastically divisible) dynamics.
  • Using the NETFD correspondence principle, we propose a structural decomposition of backflow into classical mixing and intrinsic thermo-field entanglement sectors, yielding a unified classification of overshoot and revival phenomena.
  • We explicitly demonstrate how previously obtained minimal-model phase diagrams (Figs. 23) and thermo-field entanglement results are recovered as special cases of the general TC/TCL+NETFD framework [19, 20].
2 Preliminaries: TC and TCL Master Equations
2.1 Projection operators and reduced dynamics
Let $\rho(t)$ denote the density operator of a total system, evolving under a Liouville equation
$\frac{d}{dt}\rho(t)=\mathcal{L}\rho(t),$ (1)
where $\mathcal{L}$ is the Liouville superoperator (including Hamiltonian and dissipative parts if needed). We introduce a projection superoperator $\mathcal{P}$ and its complement $\mathcal{Q}=\mathbb{I}-\mathcal{P}$. The reduced state is $\rho_{R}(t):=\mathcal{P}\rho(t)$.
The projection-operator method yields two equivalent representations of the reduced dynamics: the time-convolution (TC) equation and the time-convolutionless (TCL) equation [1, 2, 4, 3, 5, 6].
2.2 Time-convolution (TC) form
The TC equation is an integro-differential equation of the form
$\frac{d}{dt}\rho_{R}(t)=\int_{0}^{t}ds\,\mathcal{K}(t-s)\rho_{R}(s)+\eta(t),$ (2)
where $\mathcal{K}(\tau)$ is a memory kernel superoperator and $\eta(t)$ is an inhomogeneous term depending on initial correlations in the $\mathcal{Q}$-sector.
2.3 Time-convolutionless (TCL) form
The TC and TCL forms can be derived systematically in the interaction picture. The NETFD textbook explicitly distinguishes the interaction-picture TC-type equation and TCL-type equation and treats them as complementary representations of the same reduced dynamics.
The TCL equation is a time-local differential equation
$\frac{d}{dt}\rho_{R}(t)=\mathcal{G}(t)\rho_{R}(t)+\xi(t),$ (3)
where $\mathcal{G}(t)$ is the TCL generator and $\xi(t)$ is the corresponding inhomogeneous term. The two descriptions are equivalent under appropriate assumptions, and the TCL generator $\mathcal{G}(t)$ encodes memory effects via its explicit time dependence.
In what follows, we focus on the homogeneous case $\eta(t)=\xi(t)=0$, which corresponds to either factorized initial states or to situations in which the inhomogeneous terms are negligible after a short transient.
2.4 Interaction-picture TC/TCL equations in textbook form
For concreteness, we record the explicit interaction-picture TC and TCL forms as presented in the NETFD textbook. In the stochastic-Liouville derivation of quantum Langevin equations, the TCL-type equation is written as
$\frac{d}{dt}A(t)$ $=U^{-1}(t,0)\,i\left(L_{S}+PL_{I}(t)\right)e^{iL_{S}t}A$
$\quad+U^{-1}(t,0)\,QL_{I}(t)\left[1-E^{-1}(t)\right]\,iL_{I}(t)e^{iL_{S}t}A+QU^{-1}(t,0)\,b(t,0)\left[1-E^{-1}(t)\right]\,iL_{I}(t)e^{iL_{S}t}A,$ (4)
which is identified as the time-convolutionless (TCL) equation (Eq. (7.4.159) in the textbook). Starting instead from the TC representation yields the time-convolution (TC) equation
$\frac{d}{dt}A(t)=iL_{S}(t)A(t)-\int_{0}^{t}ds\,U^{-1}(s,0)\expectationvalue{L_{I}(s)L_{I}(t)}_{R}A+iL_{I}(t)\,\tilde{a}(t),$ (5)
which corresponds to Eq. (7.4.163) in the textbook.
Although these two forms appear different, they are mathematically equivalent representations of the same reduced dynamics. Their difference becomes crucial only after approximation, in which the TCL form is often preferable for identifying time-dependent effective rates and for defining phase diagrams of divisibility breaking.
3 NETFD Correspondence and Embedding Interpretation
3.1 Doubling construction
Non-equilibrium thermo field dynamics (NETFD) introduces a doubled Hilbert space
$\mathcal{H}\to\mathcal{H}\otimes\tilde{\mathcal{H}}$, (6)
[…]
3.2 Correspondence principle
A key correspondence principle in NETFD is that the master equation for the reduced density operator can be mapped to a Schrödinger-type evolution in an enlarged (doubled) space. In the notation of the NETFD textbook, the correspondence principle is summarized by the mapping
$\hat{H}_{S}X$ $\leftrightarrow[H_{S},X],$ (8)
[…]
together with the associated dynamical embedding relations (see Eqs. (8.4.112)–(8.4.115) in the NETFD formulation). This construction turns dissipative density-operator evolution into a vector evolution on an extended Hilbert space, and provides an explicit embedding sector in which information can be temporarily stored and later returned.
[…]
where $\hat{H}_{\mathrm{eff}}(t)$ is a Liouville-space effective Hamiltonian (generally non-Hermitian). This correspondence can be interpreted as an embedding of dissipative density-operator evolution into a unitary-like (or pseudo-unitary) evolution in an enlarged space [15, 14, 16, 7].
This embedding interpretation provides a structural basis for information backflow: the observable sector $\mathcal{H}$ can temporarily lose information to the tilde sector $\tilde{\mathcal{H}}$, which may later return as a revival in reduced observables.
4 Information Backflow Functional
4.1 General definition
Let $I(t)$ be a scalar information quantity associated with the reduced state $\rho_{R}(t)$. Typical examples include:
[…]
  • relative entropy $D(\rho_{R}\|\sigma)$ with respect to a reference state $\sigma$,
  • trace distance or other distinguishability measures.
We define the information backflow functional
Definition 1 (Backflow functional).
For a differentiable function $I(t)$, define
$N_{I}:=\int_{0}^{\infty}dt\,\Theta(\dot{I}(t))\,\dot{I}(t),$ (12)
where $\Theta$ is the Heaviside step function.
Thus $N_{I}$ measures the total amount of increase in $I(t)$ over the entire relaxation process. If $I(t)$ is monotone non-increasing, then $N_{I}=0$.
4.2 Interpretation
The functional $N_{I}$ does not depend on the particular microscopic mechanism of memory. It is therefore suitable as a universal diagnostic for transient overshoots and revivals. The main question is to characterize conditions under which $N_{I}=0$ must hold, and to identify structural mechanisms that allow $N_{I}>0$.
5 Main Results: Structural Conditions for Monotonicity and Backflow
This section contains the main results of the paper. The central goal is to connect information backflow to the structural properties of the reduced generator. The key advantage of working in the TCL representation is that memory effects are encoded in a time-local object.
5.1 Setting: TCL dynamics and reference states
We assume the reduced state $\rho_{R}(t)$ obeys a homogeneous TCL master equation
$\frac{d}{dt}\rho_{R}(t)=\mathcal{G}(t)\rho_{R}(t),$ (13)
where $\mathcal{G}(t)$ is a time-dependent superoperator. We also assume the existence of a stationary (or reference) state $\sigma$ satisfying
$\mathcal{G}(t)\sigma=0\qquad\text{for all }t\geq 0.$ (14)
In many dissipative models $\sigma$ is the equilibrium state.
[…]
5.2 Backflow functional
For an information observable $I(t)$ computed from $\rho_{R}(t)$, define the backflow functional
$N_{I}:=\int_{0}^{\infty}dt\,\Theta(\dot{I}(t))\,\dot{I}(t),$ (15)
which measures the total increase episodes of $I(t)$. In particular, $N_{I}=0$ if $I(t)$ is monotone non-increasing.
5.3 Theorem 1: CP-divisible TCL implies $N_{I}=0$ for relative entropy
Theorem 1 (CP-divisible TCL implies monotonicity of relative entropy).
Assume that the TCL equation (13) generates a family of dynamical maps $\Phi(t,0)$ that is completely positive and trace-preserving (CPTP) and CP-divisible: for every $t\geq s\geq 0$ there exists a CPTP map $\Phi(t,s)$ such that
$\Phi(t,0)=\Phi(t,s)\circ\Phi(s,0).$ (16)
Let $\sigma$ be an invariant state: $\Phi(t,0)\sigma=\sigma$ for all $t$. Then the quantum relative entropy is monotone non-increasing:
$D(\rho_{R}(t)\|\sigma)\leq D(\rho_{R}(s)\|\sigma),\qquad t\geq s.$ (17)
Consequently, for $I(t):=-D(\rho_{R}(t)\|\sigma)$ the backflow functional vanishes:
$N_{I}=0.$ (18)
Proof.
CP-divisibility implies $\rho_{R}(t)=\Phi(t,s)\rho_{R}(s)$ with $\Phi(t,s)$ CPTP. The data-processing inequality gives $D(\Phi(t,s)\rho\|\Phi(t,s)\sigma)\leq D(\rho\|\sigma)$. If $\sigma$ is invariant, $\Phi(t,s)\sigma=\sigma$, yielding (17). Thus $I(t)=-D(\rho_{R}(t)\|\sigma)$ is monotone non-increasing, so $\dot{I}(t)>0$ never occurs and $N_{I}=0$. ∎
5.4 Theorem 2: Time-dependent Lindblad form as a checkable sufficient condition
Theorem 2 (Time-dependent Lindblad form implies $N_{I}=0$).
Suppose the TCL generator can be written for all $t$ in the time-dependent GKSL form [11, 12]
$\mathcal{G}(t)\rho=-i[H(t),\rho]+\sum_{k}\gamma_{k}(t)\left(L_{k}(t)\rho L_{k}^{\dagger}(t)-\frac{1}{2}\{L_{k}^{\dagger}(t)L_{k}(t),\rho\}\right),$ (19)
with $\gamma_{k}(t)\geq 0$ for all $k,t$. Then the dynamics is CPTP and CP-divisible, and therefore for $I(t)=-D(\rho_{R}(t)\|\sigma)$ the backflow functional satisfies $N_{I}=0$.
Proof.
A time-local generator in GKSL form with nonnegative rates generates a CPTP divisible family of maps. The result follows from Theorem 1. ∎
5.5 Proposition 1: Classical TCL generator criterion
Proposition 1 (Classical divisibility implies monotonic KL relaxation).
If $\bm{W}(t)$ satisfies, for all $t$,
$W_{ij}(t)\geq 0\quad(i\neq j),\qquad\sum_{i}W_{ij}(t)=0,$ (20)
then the induced map is a divisible stochastic map. If $\bm{\pi}$ is a stationary distribution, then the KL divergence $D_{\mathrm{KL}}(\bm{p}(t)\|\bm{\pi})$ is monotone non-increasing, and hence $N_{I}=0$ for $I(t)=-D_{\mathrm{KL}}(\bm{p}(t)\|\bm{\pi})$.
Proof.
The stated conditions ensure that $\bm{W}(t)$ is a valid (time-dependent) generator of a stochastic evolution: the off-diagonal rates are nonnegative and column sums vanish, guaranteeing normalization preservation. Divisibility follows from the existence of a stochastic propagator $\bm{T}(t,s)$ satisfying $\bm{p}(t)=\bm{T}(t,s)\bm{p}(s)$ with $\bm{T}(t,s)$ stochastic for all $t\geq s$. Then the classical data-processing inequality implies $D_{\mathrm{KL}}(\bm{T}(t,s)\bm{p}\,\|\,\bm{T}(t,s)\bm{\pi})\leq D_{\mathrm{KL}}(\bm{p}\,\|\,\bm{\pi})$. If $\bm{\pi}$ is stationary, $\bm{T}(t,s)\bm{\pi}=\bm{\pi}$, which yields monotonicity of $D_{\mathrm{KL}}(\bm{p}(t)\|\bm{\pi})$. ∎
5.6 Theorem 3: NETFD embedding and a decomposition of backflow
Theorem 3 (Backflow decomposition via NETFD).
Assume that the extended entropy admits a decomposition
$\hat{S}(t)=S_{\mathrm{cl}}(t)+S_{\mathrm{qe}}(t),$ (21)
where $S_{\mathrm{cl}}$ is the Shannon entropy of the diagonal (classical) sector and $S_{\mathrm{qe}}$ is the intrinsic thermo-field entanglement contribution. Define the corresponding backflow measures
$N_{\mathrm{cl}}:=\int_{0}^{\infty}dt\,\Theta(\dot{S}_{\mathrm{cl}}(t))\,\dot{S}_{\mathrm{cl}}(t),\qquad N_{\mathrm{qe}}:=\int_{0}^{\infty}dt\,\Theta(\dot{S}_{\mathrm{qe}}(t))\,\dot{S}_{\mathrm{qe}}(t).$ (22)
Then the total extended-entropy backflow satisfies
$N_{\hat{S}}\leq N_{\mathrm{cl}}+N_{\mathrm{qe}},$ (23)
and the pair $(N_{\mathrm{cl}},N_{\mathrm{qe}})$ provides a model-independent classification of backflow mechanisms.
6 NETFD Decomposition of Backflow
6.1 Extended density matrix and entanglement entropy
Using the thermo-field state (7), define the extended density matrix
$|.\hat{\rho}(t):=\ket{\Psi(t)}\bra{\Psi(t)}$. (24)
Tracing out the tilde degrees of freedom yields the reduced density matrix
[…]
We define the thermo-field entanglement entropy
[…]
In many models, $\hat{S}(t)$ can be decomposed into a classical part and an intrinsic quantum-entanglement part:
$\hat{S}(t)=S_{\mathrm{cl}}(t)+S_{\mathrm{qe}}(t).$ (27)
The intrinsic part $S_{\mathrm{qe}}$ quantifies correlations not reducible to classical mixture.
6.2 Backflow decomposition
We now define a decomposition of backflow functional:
Definition 2 (Decomposed backflow).
Let $I(t)$ be a functional constructed from $\hat{S}(t)$ or related information quantities. Assume that $I(t)=I_{\mathrm{cl}}(t)+I_{\mathrm{qe}}(t)$ admits a decomposition consistent with (27). Define
$N_{I}^{(\mathrm{cl})}:=\int_{0}^{\infty}dt\,\Theta(\dot{I}_{\mathrm{cl}}(t))\,\dot{I}_{\mathrm{cl}}(t),\qquad N_{I}^{(\mathrm{qe})}:=\int_{0}^{\infty}dt\,\Theta(\dot{I}_{\mathrm{qe}}(t))\,\dot{I}_{\mathrm{qe}}(t).$ (28)
This yields a classification:
  • $N_{I}^{(\mathrm{cl})}>0$ but $N_{I}^{(\mathrm{qe})}=0$: purely classical overshoot.
  • $N_{I}^{(\mathrm{qe})}>0$: intrinsic entanglement revival (thermo-field backflow).
  • Both nonzero: hybrid backflow regime.
Remark 1.
This decomposition is meaningful because NETFD provides an explicit embedding sector (tilde degrees of freedom). The decomposition thus links information backflow to entanglement exchange between observable and auxiliary sectors.
7 Practical Algorithm for Phase-Diagram Construction
The structural results above provide a practical procedure for constructing phase diagrams of backflow from either a memory kernel (TC representation) or an effective generator (TCL representation).
Algorithm (TC $\rightarrow$ TCL $\rightarrow$ backflow classification)
  1. 1.
    Start from a reduced equation. Obtain the reduced dynamics either in TC form with memory kernel $\mathcal{K}(t)$ or in TCL form with generator $\mathcal{G}(t)$.
  2. 2.
    Compute the TCL generator. If the reduced equation is given in TC form, transform it into a TCL equation by extracting the time-local generator $\mathcal{G}(t)$ (exactly if possible, or perturbatively using standard TCL expansions).
  3. 3.
    Extract effective rates. In the quantum case, attempt to express $\mathcal{G}(t)$ in a time-dependent GKSL form (19) and identify effective rates $\gamma_{k}(t)$. In the classical case, extract the effective rate matrix $\bm{W}(t)$.
  4. 4.
    Divisibility test. If all $\gamma_{k}(t)\geq 0$ (quantum) or all off-diagonal rates are nonnegative (classical), then the dynamics is divisible and relative-entropy-type backflow is absent (Theorems 12, Proposition 1).
  5. 5.
    Compute $N_{I}$. Choose an information quantity $I(t)$ (e.g. $D(\rho\|\sigma)$, $\hat{S}(t)$, or intrinsic entropy $S_{\mathrm{qe}}(t)$) and compute $N_{I}$ by numerical integration of (15).
  6. 6.
    NETFD classification. If NETFD decomposition is available, compute $(N_{\mathrm{cl}},N_{\mathrm{qe}})$ to classify the transient as classical overshoot, intrinsic entanglement revival, or hybrid backflow.
This algorithm produces a phase diagram in the space of kernel parameters (e.g. fractional order $\alpha$, memory strength, coupling rates) and clarifies which regions correspond to divisibility breaking and to intrinsic entanglement revival.
8 Minimal Examples
Figure 1: Information backflow as temporary storage in hidden degrees of freedom and subsequent return to observables. Adapted from [20].
Figure 2: Fractional non-Markovian two-state relaxation: intrinsic thermo-field entanglement sector and phase structure. Adapted from [20].
Figure 3: Classical three-state non-Markovian relaxation: phase map of entropy/KL overshoot and amplification regimes. Adapted from [20].
8.1 Quantum two-state relaxation
Figure 4: Markovian two-state relaxation: intrinsic component $b_{\mathrm{qe}}(t)$ in the thermo-field entanglement description. Adapted from [19].
Consider a minimal two-level system with reduced density matrix dynamics described by a dissipative generator. In Markovian models, the intrinsic thermo-field entanglement component often takes the form
$b_{\mathrm{qe}}(t)=\frac{1}{4}e^{-\lambda t}\sin^{2}(\omega t),$ (29)
with decay rate $\lambda$ and oscillation frequency $\omega$. The corresponding intrinsic entanglement entropy $S_{\mathrm{qe}}(t)$ exhibits damped oscillations but no revival beyond the Markov envelope.
In non-Markovian fractional relaxation, a universal modification replaces the exponential envelope by a Mittag–Leffler function:
$b_{\mathrm{qe}}^{(\alpha)}(t)=\frac{1}{4}\left[E_{\alpha}\left(-\lambda^{\alpha}t^{\alpha}\right)\right]^{2}\sin^{2}(\omega t),$ (30)
where $0<\alpha<1$ is the fractional order and $E_{\alpha}$ is the Mittag–Leffler function [17, 18]. The deviation from exponential damping permits entanglement revivals and positive $N_{I}^{(\mathrm{qe})}$.
8.2 Classical multi-state models
For classical generalized master equations with memory kernels, one may embed the non-Markovian process into a higher-dimensional Markovian system. The TCL generator $\bm{W}(t)$ can acquire temporarily negative effective rates, corresponding to the loss of divisibility. In such cases, classical relative entropy may exhibit overshoot, yielding $N_{I}^{(\mathrm{cl})}>0$.
Fractional kernels lead again to Mittag–Leffler envelopes and can generate phase boundaries in parameter space (e.g. $\alpha$ and coupling strength) separating monotone relaxation from overshoot regimes.
9 Relation to Prior Minimal-Model Results
The present work is designed to unify and generalize two classes of analytically solvable minimal models: (i) thermo-field entanglement descriptions of two-state relaxation processes, and (ii) phase-diagram classifications of information backflow based on generalized memory kernels.
In particular, the Markovian two-state thermo-field entanglement formulation developed in [19] is recovered as a special case of the NETFD embedding framework described in Section 4. Likewise, the phase-diagram construction based on the backflow functional $N_{I}$ in [20] is obtained as an application of the structural monotonicity criteria (CP-divisibility / stochastic divisibility) in Section 5.
The new contribution of the present paper is that both results can be derived from the same TC/TCL and NETFD correspondence principles, yielding a generator-based criterion for backflow and a systematic decomposition into classical and intrinsic thermo-field contributions. This structural formulation clarifies which aspects of the phase boundaries are universal (kernel-driven) and which depend on the choice of information metric.
10 Discussion and Outlook
We proposed a structural framework that connects information backflow to the TC/TCL master equation structure and to NETFD embedding. Our main conceptual result is that backflow is absent whenever the TCL generator generates a divisible CPTP dynamics, whereas violations of divisibility provide a sufficient mechanism for backflow. NETFD provides a natural decomposition of backflow into classical and intrinsic entanglement sectors, leading to a unified classification of overshoots and revivals.
Future directions include: (i) rigorous necessary and sufficient conditions in terms of memory kernels, (ii) extension to continuous-variable Gaussian systems, and (iii) experimental diagnostics based on accessible thermodynamic observables, including entropy production bounds [13].
Appendix A Sharp decomposition in two-state NETFD parametrization
In the minimal two-state NETFD formulation, the extended reduced density matrix can be written in a canonical representation as
$\hat{\rho}_{A}(t)=\begin{pmatrix}p(t)&c(t)\\ c(t)^{\ast}&1-p(t)\end{pmatrix}$, (31)
where $p(t)$ denotes the population of one state and $c(t)$ encodes the intrinsic thermo-field sector.
Following the parametrization used in Ref. [19], we define the intrinsic contribution parameter
$b_{\mathrm{qe}}(t):=|c(t)|^{2}.$ (32)
The extended entropy can be decomposed as
$\hat{S}(t)=S_{\mathrm{cl}}(p(t))+S_{\mathrm{qe}}(b_{\mathrm{qe}}(t)),$ (33)
where $S_{\mathrm{cl}}$ is the classical mixing entropy and $S_{\mathrm{qe}}$ is a monotone function of $b_{\mathrm{qe}}$.
Taking the time derivative gives
$\dot{\hat{S}}(t)=S_{\mathrm{cl}}^{\prime}(p)\dot{p}(t)+S_{\mathrm{qe}}^{\prime}(b_{\mathrm{qe}})\dot{b}_{\mathrm{qe}}(t).$ (34)
In damped oscillatory regimes characteristic of fractional or memory-induced relaxation, both $p(t)$ and $b_{\mathrm{qe}}(t)$ are governed by the same relaxation envelope. Their positive-slope intervals therefore coincide, which yields the sharp additivity relation
$N_{\hat{S}}=N_{\mathrm{cl}}+N_{\mathrm{qe}}.$ (35)
This completes the argument underlying the sharp decomposition proposition.